OpenStudy (anonymous):

kelli adds 5 new song to her MP3 player each week after 4 weeks she has 42 songs which equation in slope intercept form represents the total number of songs she has on her MP3 player after x weeks

OpenStudy (chris):

ok so they're asking you how many songs she will have after x weeks, in slope intercept form

OpenStudy (chris):

they're giving you the slope - which is the rate of change each unit (in this case a week)

OpenStudy (chris):

that's one part of the slope intercept equation

OpenStudy (chris):

the other part is the constant

OpenStudy (chris):

from the information given, we know that "m" is "5" in the y=mx+b slope intercept equation - i.e. each week (unity on the x-axsis), 5 songes are added.

OpenStudy (chris):

so after four weeks, y (the number of songs) is 42, and we know the number of songs added each week was 5 - so really we just need the constant - the number of songs that kelli started with

OpenStudy (chris):

so y = mx +b, 42 = 5*4 +b

OpenStudy (chris):

42 = 20 + b

OpenStudy (chris):

b = 42 - 20 = 22

OpenStudy (chris):

so kellie started with 22 songs on the mp4 player (the constant), and adds 5 songs each week. the slope intercept equations is hence y= 5*x + 22

OpenStudy (chris):

does that makes sense?

OpenStudy (chris):

any particular part that doesn't make sense?